# Optimal Transport for $e/\pi^{0}$ Particle Classification in LArTPC Neutrino Experiments

### v1

@@ I Introduction
Optimal Transport (OT) is a mathematical framework that furnishes a notion of the distance between two probability distributions in terms of the minimal “work” required to rearrange one into another. In recent years OT methods have proven powerful for image processing, which suggests their relevance to data analysis in high energy physics (HEP) experiments. Indeed, OT methods have been fruitfully applied to the analysis of Large Hadron Collider (LHC) data [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], where calorimeter deposits are treated as probability distributions and OT defines a distance between events. These distances can be used for event classification and anomaly detection on their own or coupled to simple distance-based machine learning methods. The performance of these methods is often competitive with standard deep learning approaches, but more amenable to interpretation and uncertainty quantification. In this respect, OT provides a promising foundation for interpretable, physics-informed machine learning methods in HEP.
@@ II.1 Physics Landscape with LArTPC Detectors
Because of their imaging capabilities, LArTPC detectors have become the technology of choice for HEP experiments operating in accelerator-based neutrino beams. LArTPCs are able to capture images of particle interactions with millimeter position accuracy and percent-level calorimetric resolution [49, 50, 51]. This, in turn, allows for excellent particle identification (PID) which is leveraged to perform precision measurement of neutrino oscillations [52] and searches for BSM particles [53, 54, 55].
@@ II.2 Introduction to LArTPC detectors
Charged particles produced in neutrino interactions propagate through the LArTPC, ionizing and exciting the argon atoms that make up the active mass of the detector. An external electric field causes ionization electrons to drift uniformly from the bulk of the detector to the positive potential wall (anode) of the TPC. Here an array of wires measure the induced current caused by the propagating electron clouds. Charge reaches different wires at different times, depending on the exact location where the energy was deposited. Multiple overlapping wire-planes allow for the formation of different 2D views of the charge deposition, which are in turn used to reconstruct a 3D image of the interaction [60, 61]. In the GeV energy-scale range, different particle species give rise to different characteristic topologies. Particles such as protons, charged pions, or muons lose energy primarily through ionization, producing linear “track-like” signatures in the detector. Electrons and photons, on the other hand, lose energy via ionization and radiation in roughly equal amounts, leading to the formation of electromagnetic (EM) showers which appear as “fuzzy” branches and $e^{+}e^{-}$ pair cascades. Below one GeV, these EM showers are often highly segmented due to the stochastic nature of radiative energy loss. A more in-depth description of EM signatures in LArTPCs can be found in Refs. [62, 63]. Figure 1 shows an example neutrino interaction image collected by the MicroBooNE LArTPC, where different final-state particles can be seen producing topologically distinct signatures.
Figure 1: Example event display of a neutrino interaction collected with the MicroBooNE LArTPC detector. Four photon EM showers, likely from the decay of two $\pi^{0}$ mesons, can be seen branching out from the interaction vertex.
@@ III.2 Detector Simulation and Calibrations
  • The Booster Neutrino Beamline flux [66], a $\sim$1 GeV neutrino energy beamline comprised of 95% $\nu_{\mu}$, 4% $\bar{\nu}_{\mu}$, and $<1\%$ $\nu_{e}$/$\bar{\nu}_{e}$.
@@ IV.1 Current Status of LArTPC Reconstruction for $e/\pi^{0}$
Many physics results produced with LArTPC neutrino experiments leverage the Pandora, Wire-Cell, or ML tools for electron-photon separation and $\pi^{0}$ rejection [86, 87, 88, 48, 46, 84]. While these have demonstrated good performance, powerful $\pi^{0}$ rejection has come at the cost of a loss of efficiency that can negatively impact an analysis’ sensitivity. As an example, $\nu_{e}$ efficiencies for recent analyses in MicroBooNE hover around $10-40$%, with room for improvement. Furthermore, specific topologies such as isolated single-electrons with no hadronic activity [86], or more recent BSM searches for single-photons [87, 88, 48] are still heavily dominated by $\pi^{0}$ backgrounds to the point that this is the primary limitation to the analysis. Novel ways of rejecting $\pi^{0}$ backgrounds can therefore help expand the physics reach of analyses targeting single-electron or single-photon final-states. This is particularly true for methods that can overcome the reconstruction inefficiencies that often lead to not reconstructing one of the two photons from a $\pi^{0}$ decay.
Some of the ML tools developed aim to perform particle identification directly through a CNN, without the need to actually reconstruct the individual particles produced in an interaction (e.g. Ref. [78]). Such tools have been used to distinguish electrons from photons and other particles, but have not tackled the $e$/$\pi^{0}$ classification task. This paper explores OT as a new tool to tackle single-electron vs. $\pi^{0}$ classification through an image-based interpretation of the data, without the need to identify and reconstruct the two photons individually. The approach therefore is novel in that it follows a philosophy similar to that of many ML applications, which “bypass” several low-level reconstruction steps which may introduce significant inefficiencies, but does so introducing a new methodology based on OT distances, and tackling a classification problem that remains a key bottleneck in LArTPC physics analyses.
@@ V Optimal Transport Distances for LArTPC Detectors
V.1 Review of OT Distances
OT distances yield a mathematically rigorous notion of distance between probability distributions. Let $\mathcal{E}=\sum_{i=1}^{n}p_{i}\delta_{x_{i}}$ and $\mathcal{E^{\prime}}=\sum_{j=1}^{n}p^{\prime}_{j}\delta_{x^{\prime}_{j}}$ be two discrete probability distributions (which we will refer to as “events”) defined over the same underlying $d$-dimensional metric space. In other words, the event, $\mathcal{E}$, is a collection of $n$ points in $d$-dimensional space each weighted with probability $p_{i}$. OT distances elevate the pairwise distances between the points in $\mathcal{E}$ and $\mathcal{E^{\prime}}$ ($|x_{i}-x^{\prime}_{j}|$ for all possible combinations of $i,j$) to a distance between the events $\mathcal{E},\mathcal{E^{\prime}}$. Intuitively, this is done by calculating the “work” to transform $\mathcal{E}\rightarrow\mathcal{E^{\prime}}$ by moving the individual probability point masses. Moving a single probability point mass with weight $p$ from $x$ to $y$ costs an amount proportional to the distance traveled, $x-y$, and the amount of probability mass, $p$, moved. The total cost (“work”) of moving all points in $\mathcal{E}\rightarrow\mathcal{E^{\prime}}$ is the sum of these individual costs. And the transport plan which minimizes this cost is the optimal transport distance.
In this work we consider a particular kind of balanced OT distance, the 2-Wasserstein distance, which is defined as follows,
@@ V.1 Review of OT Distances
Balanced 2-Wasserstein distances are one type of a broad array of OT distances. In the next section we motivate this choice in the context of $e/\pi^{0}$ classification.
@@ V.2 $e/\pi^{0}$ Classification with OT
In order to apply OT to the task of $e/\pi^{0}$ classification, we first note that we can view the energy deposits as discrete probability distributions in 3D space. The amount of energy deposited (intensity) at a given spacepoint, up to a normalization factor, becomes the amount of probability mass. Since the shapes of these events are visually indicative of their identity, its natural to wonder whether OT distances can help in this task.
@@
We note that we also explored a different OT variant, Gromov-Wasserstein distances [89], as it is inherently insensitive to isometries (i.e. rotations and translations) of points in the ground space. Therefore, this strategy could serve as an alternative to 3D alignment (discussed in the next section). In a limited run on a single energy bin, Gromov-Wasserstein was shown to have similar performance to 3D alignment but did not see the same boost from post-processing the OT distances with interpretable machine learning strategies. Due to this lower performance as well as the high computational cost of traditional Gromov-Wasserstein algorithms, we opted to use 3D alignment instead. Results of the limited run are shown and discussed further in Appendix A.
@@ VI.2.2 Spatial Alignment
To calculate the principal axis we use Weighted Principal Component Analysis (WPCA) [91]. However, a subtlety arises when evaluating PCA in the presence of the two showers in $\pi^{0}$ events: for some events, using WPCA directly results in first principal components which are not representative of any of the shower axes. To circumvent this issue we first cluster spacepoints together if they are within 2.5 cm distance of each other and compute WPCA on the cluster with the largest number of spacepoints. The first principal component is then used to define the principal axis of the event.
@@ VI.3 Projections
Hence, inspired by event classification for jets at the LHC, we also explore using 2D projections along the principal axis component, which is analogous to the so-called jet images in colliders. Figure. 6 shows a comparison of OT plans for an $e$-$\pi^{0}$ pair (left) and two electrons (right) performed both in 3D (insets in each figure) as well as within the 2D charge projection (main panel). The particularly long distance covered by a small component of the electron in cyan along the vertical axis (inset of Fig. 6b) causes a 3D OT distance which is much larger than the 3D OT distance for the corresponding $e$-$\pi^{0}$ pair (Fig. 6a). The 2D projected charge mitigates this, leading (in this specific example) to a much smaller computed 2D OT distance for the $e$-$e$ pair compared to the $e$-$\pi^{0}$ pair.
@@ VII Methods
Optimal transport distance computations are carried out using the Python Optimal Transport (POT) library [92], which provides several generic OT solvers. In order to have a balanced sample, the same number of $\pi^{0}$ and $e$ events are chosen for each energy bin so we have $N_{p}=N_{e}$. To allow for adequate sample statistics in each bin, the bin boundaries are chosen to enforce $N_{e,p}$ to each be greater than 1,000. The bin edges are defined as [0.05, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.5, 0.9] GeV. For each bin, these $N_{p}+N_{e}$ events are divided into two equally sized training and testing sets, $N_{p,train}+N_{e,train}$ and $N_{p,test}+N_{e,test}$, respectively. Given a test event, of unknown type, the OT distance between the test event and all events in the training set is computed. The goal of each method is to assign a single number (score) which indicates the likely identity of that test event; the exact details depend on the particular strategy employed. We explore “Minimum OT”, k-Nearest Neighbors, and Support Vector Machine and describe each below.
@@ VII.1 Minimum OT: min(OT)
The first method we employ takes advantage of the observation that the OT distance between a $\pi^{0}$ and electron event will generally be larger than the distance between two electrons. The strategy employed is to designate a number of electron events as a reference, $N_{e,ref}$ and calculate the distance between a test event of unknown type and these reference events to get $N_{e,ref}$ total OT distances. We then take the minimum of these distances as our score for that test event. Namely, the score is the OT distance to the closest electron event. This approach follows from previous work on LHC data [22] which explored several methods of reduction.
@@
We reserve half of the electrons to act as a reference sample and the other half are reserved for testing; the same number of $\pi^{0}$s similarly are selected for testing. To get uncertainty estimates on our results, we further split our testing sample containing both $\pi^{0}$s and electrons into 5 randomly shuffled sets, $N^{i}_{test}$ where $i=1,...,5$. The minOT distance between each test event and the $N_{e,ref}$ electrons is computed. We can then obtain ROC curves and resulting metrics (e.g. accuracy) from various threshold cuts on min(OT) scores. We repeat this process for each $i=1,...,5$ and report the average and standard deviations of the results.
VII.2 Post-processing with Machine Learning
Interpretable ML methods can be used on OT distances to further boost the classification performance. The two methods we consider are k-nearest neighbors (kNN) and support vector machine (SVM). These methods overcome the fact that while two distributions may not be perfectly separated, OT distances for like-events will tend to cluster together, providing additional information that can be used for event classification. These methods have also been shown to boost performance in certain contexts [38, 22].

### v2

@@ I Introduction
Optimal Transport (OT) is a mathematical framework that furnishes a notion of the distance between two probability distributions in terms of the minimal “work” required to rearrange one into another. In recent years OT methods have proven powerful for image processing, which suggests their relevance to data analysis in high energy physics (HEP) experiments. Indeed, OT methods have been fruitfully applied to the analysis of Large Hadron Collider (LHC) data [83, 2, 49, 92, 95, 69, 62, 90, 91, 58, 65, 80, 79, 89, 53, 82, 100, 56, 106, 71, 63, 57, 66, 86, 85, 78, 87, 59, 1, 61, 60]. These applications have taken a variety of forms. For example, recently OT transport plans have been used to calibrate ATLAS flavour-tagging algorithms [2], to decorrelate observables [49], generate templates [92], and estimate backgrounds [95]. The majority of applications of OT to data analysis rely on calculating distances between distributions of interest. For example, OT distances have been used on Dalitz distributions to measure CP violation [69] or on the loss functions of trained VAEs to help identify anomalies [62]. However, the applications most relevant to the current work choose to represent a collider event or object (e.g. jet) as a discrete distribution over an underlying ground space [90, 91, 58, 65, 80, 79, 89, 53, 82, 100, 56, 106, 71, 63, 57, 66, 86, 85, 78, 87, 59, 1, 61, 60]. Subsequently, OT distances are calculated between these event representations [90, 91, 58, 65, 80, 79, 89, 53, 82, 100, 56, 106, 71, 63, 57, 66, 86, 85, 78, 87] or between the event representation and a constructed reference [78, 87, 59, 1, 61, 60] to define a similarity metric. Note that for a certain choice of cost-function the latter can implicitly approximate the pairwise distances between events [57, 56].
In general, OT distances between events can be used for event classification and anomaly detection on their own or coupled to simple distance-based machine learning methods. The performance of these methods is often competitive with standard deep learning approaches, but more amenable to interpretation and uncertainty quantification. In this respect, OT provides a promising foundation for interpretable, physics-informed machine learning methods in HEP.
@@ II.1 Physics Landscape with LArTPC Detectors
Because of their imaging capabilities, LArTPC detectors have become the technology of choice for HEP experiments operating in accelerator-based neutrino beams. LArTPCs are able to capture images of particle interactions with millimeter position accuracy and percent-level calorimetric resolution [12, 45, 25]. This, in turn, allows for excellent particle identification (PID) which is leveraged to perform precision measurement of neutrino oscillations [24] and searches for BSM particles [16, 22, 26]. These detectors complement other technologies employed for accelerator-based neutrino experiments, including water-Cherenkov detectors such as the next-generation Hyper-Kamiokande [6] experiment.
@@ II.2 Introduction to LArTPC detectors
Charged particles produced in neutrino interactions propagate through the LArTPC, ionizing and exciting the argon atoms that make up the active mass of the detector. An external electric field causes ionization electrons to drift uniformly from the bulk of the detector to the positive potential wall (anode) of the TPC. Wires [76, 25, 12], strips [11], or pixel-based [3, 10] charge readout sensors measure the induced current caused by the propagating electron clouds. Charge reaches these sensors at different times, depending on the exact location where the energy was deposited. In the case of wire- or strip-based detectors, multiple overlapping planes allow for the formation of different 2D views of the charge deposition, which are in turn used to reconstruct a 3D image of the interaction [40, 23]. Pixel-based detectors inherently provide 3D reconstructed charge deposits. In the GeV energy-scale range, different particle species give rise to different characteristic topologies. Particles such as protons, charged pions, or muons lose energy primarily through ionization, producing linear “track-like” signatures in the detector. Electrons and photons, on the other hand, lose energy via ionization and radiation in roughly equal amounts, leading to the formation of electromagnetic (EM) showers which appear as “fuzzy” branches and $e^{+}e^{-}$ pair cascades. Below one GeV, these EM showers are often highly segmented due to the stochastic nature of radiative energy loss. A more in-depth description of EM signatures in LArTPCs can be found in Refs. [38, 46]. Figure 1 shows an example neutrino interaction image collected by the MicroBooNE LArTPC, where different final-state particles can be seen producing topologically distinct signatures.
Figure 1: Example event display of a neutrino interaction collected with one wire plane of the MicroBooNE LArTPC detector. Four photon EM showers, likely from the decay of two $\pi^{0}$ mesons, can be seen branching out from the interaction vertex. The x-axis represents the wire position while the y-axis indicates the time taken for the electrons to drift to the wire plane. The color scale in the event display shows the measured charge deposit.
@@ III.2 Detector Simulation and Calibrations
  • The Booster Neutrino Beamline flux [105], a $\sim$1 GeV neutrino energy beamline comprised of 95% $\nu_{\mu}$, 4% $\bar{\nu}_{\mu}$, and $1\%$ $\nu_{e}$$\bar{\nu}_{e}$.
@@ IV.1 Current Status of LArTPC Reconstruction for $e/\pi^{0}$
Many physics results produced with LArTPC neutrino experiments leverage the Pandora, Wire-Cell, or ML tools for electron-photon separation and $\pi^{0}$ rejection [28, 32, 30, 33, 29, 37]. While these have demonstrated good performance, powerful $\pi^{0}$ rejection has come at the cost of a loss of efficiency that can negatively impact an analysis’ sensitivity. As an example, $\nu_{e}$ efficiencies for recent analyses in MicroBooNE hover around 104010-40%, with room for improvement. Furthermore, specific topologies such as isolated single-electrons with no hadronic activity [28], or more recent BSM searches for single-photons [32, 30, 33] are still heavily dominated by $\pi^{0}$ backgrounds to the point that this is the primary limitation to the analysis. Novel ways of rejecting $\pi^{0}$ backgrounds can therefore help expand the physics reach of analyses targeting single-electron or single-photon final-states. This is particularly true for methods that can overcome the reconstruction inefficiencies that often lead to not reconstructing one of the two photons from a $\pi^{0}$ decay.
[…]
@@ V Optimal Transport Distances for LArTPC Detectors
V.1 Review of Discrete OT Distances
OT distances yield a mathematically rigorous notion of distance between probability distributions. Let $\mathcal{E}=\sum_{i=1}^{n}p_{i}\delta_{x_{i}}$ and $\mathcal{E^{\prime}}=\sum_{j=1}^{n^{\prime}}p^{\prime}_{j}\delta_{x^{\prime}_{j}}$ be two discrete probability distributions (which we will refer to as “events”) defined over the same underlying $d$-dimensional metric space. [^1] Note that $n$ need not be equal to $n^{\prime}$. However, for balanced OT distances the total probability mass must be the same in each event i.e. $\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{n^{\prime}}p^{\prime}_{j}$ In other words, the event, $\mathcal{E}$, is a collection of $n$ points in $d$-dimensional space each weighted with probability $p_{i}$. OT distances elevate the pairwise distances between the points in $\mathcal{E}$ and $\mathcal{E^{\prime}}$ ($|x_{i}-x^{\prime}_{j}|$ for all possible combinations of $i,j$) to a distance between the events $\mathcal{E},\mathcal{E^{\prime}}$. Intuitively, this is done by calculating the “work” to transform $\mathcal{E}\rightarrow\mathcal{E^{\prime}}$ by moving the individual probability point masses. Moving a single probability point mass with weight $p$ from $x$ to $y$ costs an amount proportional to the distance traveled, xyx-y, and the amount of probability mass, $p$, moved. The total cost (“work”) of moving all points in $\mathcal{E}\rightarrow\mathcal{E^{\prime}}$ is the sum of these individual costs. The optimal transport distance is the minimum cost incurred over all possible transport plans.
In this work we consider a particular kind of discrete, balanced OT distance, the 2-Wasserstein distance [57], which is defined as follows,
@@ V.1 Review of Discrete OT Distances
In this work we focus on calculating discrete OT distances between events. [^2] We note that there has been interesting work on semi-discrete [59, 1, 61, 60] alternatives in which events are represented as a discrete distribution but distances are calculated to a continuous manifold. The choice of manifold can help accentuate features in the events. However, in practice, this continuous manifold must be approximated by a discrete distribution. This is a natural choice since the energy deposits in an event can naturally be expressed as a discrete probability distribution over the detector volume. Furthermore, balanced 2-Wasserstein distances are one type of a broad array of discrete OT distances. In the next section we motivate this choice in the context of $e/\pi^{0}$ classification.
@@ V.2 $e/\pi^{0}$ Classification with OT
In order to apply OT to the task of $e/\pi^{0}$ classification, we first note that we can view the energy deposits as discrete probability distributions in 3D space. The amount of energy deposited (intensity) at a given spacepoint, up to a normalization factor, becomes the amount of probability mass. Since the shapes of these events are visually indicative of their identity, it’s natural to wonder whether OT distances can help in this task.
@@
We note that we also explored a different OT variant, Gromov-Wasserstein distances [Mémoli_2011], as it is inherently insensitive to isometries (i.e. rotations and translations) of points in the ground space. Therefore, this strategy could serve as an alternative to 3D alignment (discussed in the next section). In a limited run on a single energy bin, Gromov-Wasserstein was shown to have similar performance to 3D alignment but did not see the same boost from post-processing the OT distances with interpretable machine learning strategies. Due to this lower performance as well as the high computational cost of traditional Gromov-Wasserstein algorithms, we opted to use 3D alignment instead. Results of the limited run are shown and discussed further in Appendix A.
@@ VI.2.2 Spatial Alignment
To calculate the principal axis we use Weighted Principal Component Analysis (WPCA) [70]. However, a subtlety arises when evaluating WPCA in the presence of the two showers in $\pi^{0}$ events: for some events, using WPCA directly results in first principal components which are not representative of any of the shower axes. To circumvent this issue we first cluster spacepoints together if they are within 2.5 cm distance of each other and compute WPCA on the cluster with the largest number of spacepoints. The first principal component is then used to define the principal axis of the event.
@@ VI.3 Projections
Hence, inspired by event classification for jets at the LHC, we also explore using 2D projections along the principal axis component, which is analogous to the so-called jet images in colliders. The primary similarity here is in the construction of a physically relevant axis along which to take the projection such that angular deviations carry physical meaning. Figure. 6 shows a comparison of OT plans for an $e$-$\pi^{0}$ pair (left) and two electrons (right) performed both in 3D (insets in each figure) as well as within the 2D charge projection (main panel). The particularly long distance covered by a small component of the electron in cyan along the vertical axis (inset of Fig. 6b) causes a 3D OT distance which is much larger than the 3D OT distance for the corresponding $e$-$\pi^{0}$ pair (Fig. 6a). The 2D projected charge mitigates this, leading (in this specific example) to a much smaller computed 2D OT distance for the $e$-$e$ pair compared to the $e$-$\pi^{0}$ pair.
@@ VII Methods
Optimal transport distance computations are carried out using the Python Optimal Transport (POT) library [77], which provides several generic OT solvers. In order to have a balanced sample, the same number of $\pi^{0}$ and $e$ events are chosen for each energy bin so we have $N_{\pi^{0}}=N_{e}$. To allow for adequate sample statistics in each bin, the bin boundaries are chosen to enforce $N_{e,p}$ to each be greater than 1,000. The bin edges are defined as [0.05, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.5, 0.9] GeV. For each bin, these $N_{\pi^{0}}+N_{e}$ events are divided into two equally sized training and testing sets, $N_{\pi^{0},train}+N_{e,train}$ and $N_{\pi^{0},test}+N_{e,test}$, respectively.
Given a test event, of unknown type, each OT method must assign a single number (score) which indicates the likely identity of that test event. The exact details of how the score is derived depend on the particular strategy employed. We explore “Minimum OT”, k-Nearest Neighbors, and Support Vector Machine and describe each below.
@@ VII.1 Minimum OT: min(OT)
The first method we employ takes advantage of the fact that the OT distance between like-distributions will be smaller than the distance between unlike-distributions. Since we expect physically similar events to have similar detector signatures, we can use this property to our advantage. In this case, the OT distance between a $\pi^{0}$ event and electron event will generally be larger than the distance between two electron events.
In the min(OT) method, the strategy employed to identify $\pi^{0}$ events is to first designate a number of electron events from the training set as a reference, $\{\mathcal{E}_{e_{i}}\mid i=1,...,N_{e,ref}\}$. We can then compare an event of an unknown type, $\mathcal{T}$, to this reference population. Namely, for each event $\mathcal{T}$ we can calculate $N_{e,ref}$ OT distances $\{{\rm OT}(\mathcal{T},\mathcal{E}_{e_{i}})~\mid~i=1,...,N_{e,ref}\}$. Electron events, $\mathcal{T}_{e}$, will tend to have a smaller OT distance than $\pi^{0}$ events, $\mathcal{T}_{\pi^{0}}$ i.e. the set $\{{\rm OT}(\mathcal{T}_{e},\mathcal{E}_{e_{i}})\}$ will generally be closer to 0 than $\{{\rm OT}(\mathcal{T}_{\pi^{0}},\mathcal{E}_{e_{i}})\}$. The task now is to turn $\{{\rm OT}(\mathcal{T},\mathcal{E}_{e_{i}})~\mid~i=1,...,N_{e,ref}\}$ into a single score for each test event $\mathcal{T}$.
We choose to take the minimum of these distances as our score for that test event. Namely, the score for an event $\mathcal{T}$ is the OT distance to the closest electron event in the reference sample. This approach follows from previous work on LHC data [65] which explored several methods of reduction. We also note that there have been several works in the context of particle colliders which have argued that the notion of smallest OT distance is a valid discriminating observable [87, 78, 65]. Ref. [87] even found that in some cases traditional physics observables correspond to the closest OT distance to a manifold of events.
@@
We reserve half of the electron events to act as a reference sample and the other half are reserved for testing; the same number of $\pi^{0}$ events similarly are selected for testing. To get uncertainty estimates on our results, we further split our testing sample containing both $\pi^{0}$s and electrons into 5 randomly shuffled sets each of size $N_{test}$. The min(OT) distance between each test event and the $N_{e,ref}$ electrons is computed. We can then obtain ROC curves and resulting metrics (e.g. accuracy) from various threshold cuts on min(OT) scores. We repeat this process for each of the 5 test sets and report the average and standard deviations of the results.
VII.2 Post-processing pairwise OT distances with Machine Learning
Alternatively, interpretable ML methods can be used on pairwise OT distances to further boost the classification performance. The two methods we consider are k-nearest neighbors (kNN) and support vector machine (SVM). These methods overcome the fact that while two distributions may not be perfectly separated, OT distances for like-events will tend to cluster together, providing additional information that can be used for event classification. These methods have also been shown to boost performance in certain contexts [57, 65].
In contrast, the min(OT) method relies on the distribution of $\min\{{\rm OT}(\mathcal{T}_{\pi^{0}},\mathcal{E}_{e_{1}}),...,{\rm OT}(\mathcal{T}_{\pi^{0}},\mathcal{E}_{e_{N_{e,ref}}})\}$ being sufficiently separated from the distribution of $\min\{{\rm OT}(\mathcal{T}_{e},\mathcal{E}_{e_{1}}),...,{\rm OT}(\mathcal{T}_{e},\mathcal{E}_{e_{N_{e,ref}}})\}$. While this allows you to operate in an unsupervised manner, you risk sacrificing additional separation power that supervised post-processing (kNN and SVM) can yield. However, these methods now require both electron and $\pi^{0}$ events for training. Namely, the inputs to these models are the full set of pairwise OT distances, $\{{\rm OT}(\mathcal{E}_{\pi^{0}},\mathcal{E}_{e})\}$, for all $e,\pi^{0}$ training events.